OpenStudy (anonymous):

How many liters of fluorine gas can react with 32.0 grams of sodium metal at standard temperature and pressure? Show all of the work used to find your answer. 2Na + F2 yields 2NaF

OpenStudy (anonymous):

I have: 32 g NaF x \[\frac{ 1 mol NAF }{ 41.98 g NaF}\] x\[\frac{ 1 mol F _{2} }{ 2 mol NaF }\] x \[\frac{ 1 L F _{2} }{ 1 mol F _{2} }\] = .381

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

@gabylovesu

OpenStudy (anonymous):

@karatechopper

OpenStudy (anonymous):

@mathstudent55

Vocaloid (vocaloid):

@FutureVet one small mistake the last step should be 22.4L F2/1 mol F2, because 1 mol of gas at standard temperature and pressure occupies 22.4L

Vocaloid (vocaloid):

other than that, everything else looks ok

OpenStudy (anonymous):

Thanks I realized that before I Submitted it.